### v3

# Pauli Check Sandwiching for Quantum Characterization and Error Mitigation during Runtime

Abstract
This work presents a novel quantum system characterization and error mitigation framework that applies Pauli check sandwiching (PCS). We motivate our work with prior art in software optimizations for quantum programs like noise-adaptive mapping and multi-programming, and we introduce the concept of PCS while emphasizing design considerations for its practical use. We show that by carefully embedding Pauli checks within a target application (i.e. a quantum circuit), we can learn quantum system noise profiles. Further, PCS combined with multi-programming unlocks non-trivial fidelity improvements.
I Introduction and Motivation
While the quantum compute stack has improved, existing quantum system noise still varies cross-chip (Fig. 1) and hovers above error correction thresholds. Much work has been dedicated to improving quantum program success. For example, noise-adaptive mapping is a state-of-art quantum circuit optimization technique used to place and route quantum programs on quantum hardware  [5]. Noise-aware schemes, however, depend on accurate calibration data from the quantum computer (QC) provider or from benchmarking. Even in the case that a snapshot of machine properties were available, QC noise is observed to fluctuate over time in ways that are challenging to capture with analytical models.
In our work, we seek alternative methods to determine the noise profile of a QC when its properties are unknown. It is desirable to develop characterization techniques with minimal overheads in terms of quantum processor unit (QPU) runtime. We are inspired by prior art in multi-programming  [1] and ensemble distributions [6] to explore how parallelization during runtime can be employed for 1) learning quantum noise / hardware characterization, 2) fidelity improvements in program outcomes, and 3) reductions in runtime latency. This work presents a novel framework that combines the power of multi-programming with Pauli check sandwiching (PCS) for QC characterization and error mitigation. Multi-programming is a technique for parallelized quantum circuit execution while PCS is an error mitigation strategy that utilizes post-selection to reduce errors in quantum circuit outcomes. Our techniques intelligently embed PCS into quantum applications. These programs are then simultaneously executed across a QPU for noise profile characterization. Resulting characterization data adaptively mitigates noisy outcomes within individual thread outcomes. Knowledge of the QPU’s best regions creates a weighted ensemble result with non-trivial gain, found to be up to a $25\%$ fidelity improvement in our reported results.
Figure 1: Example QPU with 20 regions (10 qubits/region) of unique 1q error rates, $p\in[0.0005,0.01]$, and 2q errors, $2p$. PCS-based characterization with multi-programmed circuits learns the QPU’s unknown noise properties.
II Pauli Check Sandwiching
We rely on a technique called ‘Pauli Check Sandwiching’ (PCS) to detect and mitigate errors [2, 4, 3]. In our work, we also find that PCS is a powerful tool for guiding QC characterization. As seen in Fig. 2, PCS surrounds a payload circuit, $U$, with controlled Pauli operator checks. Errors on $U$ can be detected on an ancilla through phase kickback. It is important that the relationship $R_{1}UL_{1}=U$ holds so the introduced checks do not disturb the original payload circuit semantics.
Recognizing that all errors can be decomposed into Pauli operators is a key concept in understanding PCS. By cleverly constructing Pauli sandwich checks dependent on $U$, anti-commutativity relations between the errors and checks can reveal errors in the ancilla qubits. Measuring 1 in any ancillas indicates that phase kickback occurred, meaning anti-commutation between the Paulis, and thus, a present error. As part of the PCS protocol, we discard this error-corrupted shot, leading to an increase in final distribution fidelity of $U$.
Practical application of PCS requires consideration of tradeoffs. First, determining the Pauli unitary used in the check increases in complexity with the circuit. Second, Pauli check logic must be decomposed into the supported gate set of the QC. Third, the overhead of adding PCS into the circuit must not outweigh the corrective benefit in terms of gate error. We consider these constraints in our solution.
Figure 2: General PCS circuit layout. The red unitaries represent the Pauli checks that sandwich the main payload circuit. Measurement of the ancillas provide detection of errors that occurred in the payload circuit.
III QC Characterization and Error Mitigation with PCS
With the PCS protocol, we can 1) learn the noise profile of the QC and 2) increase circuit fidelity by eliminating the states with errors. First, we utilize single-qubit checks only on edge qubits to minimize PCS gate count overhead. We then maximally utilize hardware via multi-programming; a QC with $n$ physical qubits holds $threads=\lfloor n/(q_{algorithm}+q_{ancilla})\rfloor$ instances of a circuit running in parallel. The value of the checks is used during post processing to filter errored outcomes from each of the threads, creating vectors of error-mitigated counts, $c_{i}$. The percentage of discarded states, $d_{i}$, gives insight into the underlying noise profile of the QC when they are compared. Each local $d_{i}$ is then used as a weight that scales $c_{i}$ to create $s_{i}=c_{i}*\frac{min(d)}{d_{i}}$. A cumulative distribution, $S$, is created from the sum of the $s_{i}$ vectors from the PCS-protected local threads, $S=\sum s_{i}$.
IV Experimental Results
We simulated the fidelity rates of PCS in both a GHZ / sensing circuit and Toffoli circuit to concretely measure the improvement. We expand the Fig. 1 model QC and study a system with 60 10-qubit regions where single-qubit depolarizing noise $p\in[0.0005,0.03]$ and two-qubit gates are $2p$. We find that PCS not only mitigates error in individual circuit distributions – check data can also be used to intelligently combine the results of multiple threads cross-QPU, demonstrating measurable noise reductions in final outcomes. Fig. 3 shows our results, detailing a multiple percentage point increase in fidelity when multi-programmed applications are protected with PCS and a cumulative distribution is created from local results weighted by check data. Even with only the edge qubits protected, a measurable benefit results. Fig. 4 documents the relationship between discarded shots and error rate, illustrating PCS’s potential for QC error characterization. We believe this suggests that PCS has additional promise to guide efficient qubit mapping.
Figure 3: Fidelity of circuits with no PCS (base) vs. PCS protection. 10,000 shots were used for both circuits in each simulator region (60 total), and fidelity improvement of the PCS weighted ensemble was determined to be $25\%$ (left) and $16.25\%$ (right) as compared to a naive ensemble.
Figure 4: Shots removed / total shots as error rates increase. 10,000 shots on a GHZ circuit were executed for every error rate (60 total).
V Acknowledgments
JL, AG, and ZHS acknowledge support by the Q-NEXT Center. NH was partially supported by NSF CCF-2119069.
The submitted manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory (“Argonne”). Argonne, a U.S. Department of Energy Office of Science laboratory, is operated under Contract No. DE-AC02-06CH11357. The U.S. Government retains for itself, and others acting on its behalf, a paid-up nonexclusive, irrevocable worldwide license in said article to reproduce, prepare derivative works, distribute copies to the public, and perform publicly and display publicly, by or on behalf of the Government. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan. http://energy.gov/downloads/doe-public-access-plan.